# Robust Stability Analysis of Positive Lur’e System with Neural Network Feedback

### v1

@@ Abstract
This paper investigates the robustness of the Lur’e problem under positivity constraints, drawing on results from the positive Aizerman conjecture and the robustness properties of Metzler matrices. Specifically, we consider a control system of Lur’e type in which not only the linear part includes parametric uncertainty but also the nonlinear sector bound is unknown. We investigate tools from positive linear systems to effectively solve the problems in complicated and uncertain nonlinear systems. By leveraging the positivity characteristic of the system, we derive an explicit formula for the stability radius of Lur’e systems. Furthermore, we extend our analysis to systems with neural network (NN) feedback loops. Building on this approach, we also propose a refinement method for sector bounds of feedforward neural networks (FFNNs). This study introduces a scalable and efficient approach for robustness analysis of both Lur’e and NN-controlled systems. Finally, the proposed results are supported by illustrative examples.
@@ I Introduction
In recent years, the rapid advancement of NNs has propelled their integration into various autonomous systems, raising critical concerns regarding their stability and robustness in real-world applications [1]. As these networks increasingly become embedded in safety-critical domains, it is imperative to ensure their reliable performance under uncertain conditions. Researchers strived to address this issue by adopting various concepts from mathematics and control theory. Surveys such as [2] and [3] explain some of the efforts. However, a very straightforward yet resourceful tool has been overlooked in these efforts—the Lur’e system. This study leverages the Lur’e system framework to address the robustness issue of NN-controlled systems.
@@
Lur’e systems provide a versatile model for representing systems with NN feedback loops, enabling the use of numerous analytical tools to assess system stability. They also offer a suitable framework for incorporating uncertainty in nonlinear systems. In this paper, we explore the robustness of Lur’e systems under structured uncertainty with positivity constraints on the system. Later, we apply this novel tool to uncertain NN-controlled systems, addressing the current safety issue of the real-world applications.
@@
Lur’e systems have long served as a foundation for stability analysis of nonlinear systems, underpinning fundamental methods such as the Lur’e-Postnikov criterion, Aizerman and Kalman conjectures, the circle criterion, and their extensions. However, these classical approaches have limitations when dealing with system uncertainties, leaving room for further exploration. The existing literature on robustness analysis of Lur’e systems with structured uncertainties, which is the focus of our study, spans various approaches. For instance, studies such as [4, 5, 6] employ different methods, including Popov criterion, $H_{\infty}$ control, and Lyapunov theory, respectively. Additionally, some works, like [7], explore more complex perturbation structures. There are recent works such as [8] that implicitly studied robust stability of Lur’e systems using Integral Quadratic Constraints (IQCs) and implemented it on an NN-controlled system. However, their proposed method suffers from high computational complexity when it comes to large systems. In contrast, this paper introduces a straightforward and scalable approach to robustness analysis of Lur’e systems by utilizing positivity constraints on the system, offering a computationally efficient alternative to existing methods.
@@
It is worth noting that positive-constrained systems are widely used in modeling numerous real-world systems such as engineering, biological, economical, population, and other systems. The robustness of linear positive systems has been extensively studied, with significant contributions from researchers such as Farina and Rinaldi [9]. The positivity constraint offers a significant advantage in the stability and robustness analysis of linear systems [10]; however, its integration into nonlinear Lur’e systems has not been extensively explored. There is some available literature, such as [11] in which the authors demonstrated that the Aizerman conjecture holds for a class of positive Lur’e systems. We aim to further explore this constraint and its advantages in the analysis of Lur’e systems. By leveraging this constraint, we can employ a fundamental linear system analysis tool to assess the robustness of complex nonlinear Lur’e systems with structured uncertainty. A key novelty of this paper is demonstrating how a foundational linear tool can effectively enhance the robustness analysis of a complex nonlinear system.
In this paper, we derive an explicit formula for the stability radius of Lur’e systems under a comprehensive class of structured uncertainties. Compared to previous studies with similar uncertainty structures, such as [12, 13], our approach offers a more streamlined and computationally efficient method for robust stability analysis.
@@
Lur’e systems have also been employed in the safety analysis of NNs. For instance, studies such as [14, 8, 15] implicitly utilize Lur’e systems alongside the S-procedure to define the forward reachable set of a NN and ensure the stability of NN-controlled systems. Most recently, a notable study [16] introduced a simple, scalable method for stability assurance in NN-controlled systems using a Lur’e framework. The study from [17] utilizes Lur’e system with reinforcement learning subject to time delays. Also, [18] investigated the stability of feedback systems with ReLU NNs. However, some of these approaches either lack a formal robustness measure or rely on solving computationally expensive Linear Matrix Inequalities (LMIs). To address this gap, our study proposes a straightforward and scalable method for calculating the stability radius of systems with NN feedback loops.
@@
In addition to robustness analysis, this study addresses sector bounds for FFNNs. Sector bounds are a classical tool in control theory, used to ensure stability in systems with nonlinearities by defining constraints within which the system’s nonlinearity must operate. NN bounds are also one of the leading topics in the Machine Learning community. There are extensive recent studies on NN bounding methods and refinement of the existing bounds. Examples are the evolution of bounding methods like CROWN [19, 20], BnB [21], IBP [22], and sector bounds [16]. In this study, we build upon our results in Lur’e systems and propose a new simple method for refining sector bounds specifically for FFNN-controlled systems.
@@ I-A Notation
The set of real numbers, $n$-dimensional real vectors, and $m\times n$-dimensional real matrices are denoted by $\mathbb{R}$, $\mathbb{R}^{n}$, and $\mathbb{R}^{m\times n}$, respectively. Additionally, $\mathbb{C}$ is set of complex numbers. The standard comparison operators $<$, $\geq$, $\leq$, and $>$ are applied elementwise to vectors and matrices. $\mathbb{R}_{+}$ denotes the set of nonnegative real numbers. Moreover, $\mathbb{R}_{+}^{n}:=\{\nu\in\mathbb{R}^{n}:\nu\geq 0\}$ represents the set of vectors with real nonnegative elements. Similarly, $\mathbb{R}_{+}^{n\times m}$ denotes the set of real matrices with nonnegative elements. A Metzler matrix is characterized by having nonnegative off-diagonal entries. We show the set of Metzler matrices by $\mathbb{M^{+}}$. An M-matrix is defined as a matrix whose off-diagonal elements are non-positive and whose eigenvalues all have non-negative real parts. Vectors of zeros and the identity matrix of size $n$ are denoted by $\mathbf{0}_{n}$ and $\mathbf{I}_{n}$, respectively. For a given matrix $A$, the order $|A|_{\text{abs}}$ means elementwise absolute value of matrix $A$, and $\|A\|$ denotes a matrix norm of choice.
@@ II Proposed Theoretical Framework
In this section, we establish the theoretical frameworks that underpin our study, including the positive Lur’e system, positive Aizerman’s conjecture, and the stability radii of nonnegative and Metzler matrices.
@@ II-A Positive Lur’e Systems
Consider a linear time invariant (LTI) system of the form:
$\small\dot{x}(t)=Ax(t)+Bu(t),\quad y(t)=Cx(t),$ (1)
@@
The system described by (1) is positive if and only if the matrix $A$ is a Metzler matrix, $B\in\mathbb{R}_{+}^{n\times m}$, and $C\in\mathbb{R}_{+}^{p\times n}$.
@@
A positive system of the form (1) is asymptotically stable if there exists a positive vector $v$ such that $Av<0$.
Now, we officially define positive Lur’e system. Consider a system formed by the interconnection of the linear system (1) and a static nonlinear feedback, represented as $u=\Phi(Cx,.)$. Fig. 1 depicts a schematic of this system. The overall dynamics of the system can be described by the following equation:
$\small\dot{x}=Ax(t)+B\Phi(Cx,t).$ (2)
@@
As part of the definition of Lur’e systems, it is necessary to establish sector bounds for the nonlinear component of the system. In the context of Multi Input Multi Output (MIMO) systems, the multivariable nature of $\Phi$ necessitates a generalization of the classical sector bound definition used in Single Input Single Output (SISO) Lur’e systems. For positive systems, the sector bound for the multivariable function $\Phi$ is best defined through componentwise inequalities. The function $\Phi$ is said to be sector bounded in $[\Sigma_{1},\Sigma_{2}]$ if the following condition holds:
@@
where $\Sigma_{1},\Sigma_{2}\in\mathbb{R}^{m\times p}$, and $\Sigma_{1}\leq\Sigma_{2}$. At this stage, the MIMO Positive Lur’e system is fully defined. With this foundation established, we can introduce the positive Aizerman’s conjecture in the next subsection.
II-B Positive Aizerman’s Conjecture
Aizerman’s conjecture is one of the absolute stability analysis methods of general Lur’e systems. The classical framework of Aizerman’s conjecture applies to SISO Lur’e systems. It is important to acknowledge that counterexamples to general Aizerman’s conjecture do exist. However, under more stringent conditions, such as system positivity, the conjecture remains valid [11]. Moreover, Aizerman’s conjecture can be extended to MIMO systems, offering substantial implications for system design. This conjecture provides a simplified criterion for predicting the stability of nonlinear systems by analyzing their linear approximations. It can thus serve as the foundation for creating scalable and efficient stability analysis techniques, particularly in systems with complex nonlinearity, such as NN-controlled systems.
The following theorem and remark is presented in [11], confirming the validity of Aizerman’s conjecture for MIMO positive systems.
@@ II-B Positive Aizerman’s Conjecture
Consider the Lur’e system described by (2) where $B\geq 0$, $C\geq 0$, and the matrices $\Sigma_{1}$ and $\Sigma_{2}$ satisfy $\Sigma_{1}\leq\Sigma_{2}$ with appropriate dimensions. If the matrix $A+B\Sigma_{1}C$ is Metzler and $A+B\Sigma_{2}C$ is Hurwitz, then for any nonlinearity $\Phi$ within the sector $[\Sigma_{1},\Sigma_{2}]$, as defined by (3), the Lur’e system (2) is globally exponentially stable.
@@
The Aizerman conjecture can be interpreted as a stability radius analysis of a linear system $\dot{x}=Ax$ subjected to an additive, structured perturbation given by:
@@
The Positive Aizerman conjecture asserts that if a system remains stable under all perturbations of a linear form $\{B\Sigma C:\quad\forall\Sigma\in[\Sigma_{1},\Sigma_{2}]\}$, then the system will also be stable for all corresponding nonlinear perturbations of the form $\{B\Phi(Cx):\forall\Phi\in[\Sigma_{1},\Sigma_{2}]\}$. In terms of stability radius, Positive Aizerman conjecture suggests that the ”real static nonlinear stability radius,” denoted $r_{\mathbb{R},\Phi}$, is equal to the ”real linear stability radius,” $r_{\mathbb{R}}$. It also implies that the stability radius of a Lur’e system can be calculated by the stability radius of its linearizations.
@@ II-C Stability Radius of Positive Systems
where $A\in\mathbb{R}^{n\times n}$ is the system’s nominal matrix. $\Delta\in\mathbb{R}^{n\times n}$ represents a perturbation matrix and $\|\Delta\|$ is a norm representing the size of the perturbation.
Suppose the following LTI system with structured perturbation:
@@
We focus on the case where $A$ is Metzler and stable. The stability assumption is included solely to guarantee the existence of the stability radius. Also, $D\geq 0$ and $E\geq 0$ are typical assumptions for systems with Metzler matrices, and simplify the stability analysis.
The following theorem provides an explicit formula for the stability radius of Metzler matrices. For this class of matrices, it has been recognized that the real and complex stability radii are coinciding and given by a direct formula. The following theorem states this result.
@@
[…]
@@
Furthermore, if $\Delta$ is defined by the set $\Delta=\{S\odot\Delta:S_{ij}\geq 0\}$ with $\|\Delta\|=\max\{|\delta_{ij}|:\delta_{ij}\neq 0\}$ where $[S\odot\Delta]_{ij}=S_{ij}\delta_{ij}$ represents the Schur product, then
@@ III Stability Radius of Positive Lur’e Systems
In this section, we formally define our problem and introduce our main result. We demonstrate and prove how we can utilize characteristics of Metzler matrices in linear systems for handling problems of nonlinear systems. Specifically, we study the stability radius of nonlinear Lur’e systems with structured perturbations under positivity constraints.
Consider the positive Lur’e system described by (2), where the nonlinearity $\Phi(.)$ satisfies the sector condition (3) in $[\Sigma_{1},\Sigma_{2}]$, with $\Sigma_{1},\Sigma_{2}\in\mathbb{R}^{m\times p}$ and $\Sigma_{1}\leq\Sigma_{2}$. We are interested in the robustness of this system to structured perturbations of the form:
@@
$r_{\mathbb{R}}(\Phi)$ $=\inf\{\|\Delta\|:\left(A+D\Delta E\right)x(t)$ (9)
$+B\Phi(Cx(t))\text{ is not globally assymptotically stable.}\}$. }
@@
Let $P,Q\in\mathbb{R}^{n\times n}$ be Metzler and Hurwitz stable matrices. Consider structured perturbations of the form
@@
$r_{\mathbb{R}}(P)\;\leq\;r_{\mathbb{R}}(Q),$ (10)
[…]
@@
Since $P$ is Metzler and Hurwitz, $-P$ has nonpositive off-diagonals and eigenvalues with positive real parts. This implies that $-P$ is an M-Matrix. The same applies to $-Q$.
Given $P\geq Q$, we have $-P\leq-Q$. Therefore, Based on a classical property of nonsingular M-matrices
@@
[…]
@@
Taking reciprocals and removing the negative sign inside the norm gives the inequality for the stability radii:
@@
We can also explain Lemma 3 in different words. It is well known that the interval positive systems $[Q,P]$ are Hurwitz stable if and only if $P$ is Hurwitz stable. This has been established in the theorem $4$ of reference [24] for positive discrete time systems, which is equally applicable for continuous-time positive systems as it is shown in [25]. Now, without loss of generality, assume $D=E=I$. Since $P>Q$ and $P$ is asymptotically stable, it is closest to instability due to continuity of stable eigenvalues. Thus, the distance of $P$ from the set of singular matrices is given by $\|P^{-1}\|^{-1}$, which is the smallest singular value of $P$ with respect to $2-$norm. In other words, $r_{\mathbb{C}}=r_{\mathbb{R}}=\sigma_{\min}=\|P^{-1}\|^{-1},$ where $\sigma$ denotes singular value of matrix $P$.
@@
Consider a Lur’e system described in (8). Given that $A+B\Sigma_{1}C$ is Metzler and $A+B\Sigma_{2}C$ is Hurwitz, the stability radius of positive Lur’e system $r_{\mathbb{R}}(\Phi)$ with respect to the structured perturbation $D\Delta E$ is given by:
@@
Based on positive Aizerman’s conjecture and remark 1, the stability radius of the nonlinear Lur’e system can be measured by the stability radius of all the linear systems of the form $A+B\Sigma C,\Sigma\in[\Sigma_{1},\Sigma_{2}]$ where $\Phi$ is sector bounded in the same interval. In other words,
$r_{\mathbb{R}}(\Phi)=\min_{\Sigma\in[\Sigma_{1},\Sigma_{2}]}r_{\mathbb{R}}(A+B\Sigma C).$ (12)
@@
$=r_{\mathbb{R}}(A+B\Sigma_{2}C)$ (13)
Now (12) and (13) result in
@@
Since the nonlinear function $\Phi$ lies within the sector $[\Sigma_{1},\Sigma_{2}]$, and $A+B\Sigma_{2}C$ is Hurwitz, based on the implications of Remark 1, the same stability radius applies to the Lur’e system. ∎
It should be pointed out that the above theorem can also be stated in terms of other uncertainty structures, such as the one stated in Theorem 2 with the aid of Eq. (7).

### v3

@@ Abstract
This paper investigates the robustness of the Lur’e problem under positivity constraints, drawing on results from the positive Aizerman conjecture and robustness properties of Metzler matrices. Specifically, we consider a control system of Lur’e type in which not only the linear part includes parametric uncertainty but also the nonlinear sector bound is unknown. We investigate tools from positive linear systems to effectively solve the problems in complicated and uncertain nonlinear systems. By leveraging the positivity characteristic of the system, we derive an explicit formula for the stability radius of Lur’e systems. Furthermore, we extend our analysis to systems with neural network (NN) feedback loops. Building on this approach, we also propose a refinement method for sector bounds of NNs. This study introduces a scalable and efficient approach for robustness analysis of both Lur’e and NN-controlled systems. Finally, the proposed results are supported by illustrative examples.
@@ I Introduction
In recent years, the rapid advancement of NNs has driven their integration into various safety critical autonomous systems, raising critical concerns about their stability and robustness in real-world applications [1]. Therefore, ensuring their reliable performance under uncertainty is imperative. Researchers strived to address this issue by adopting various concepts from mathematics and control theory. Surveys such as [2] and [3] explain some of the efforts. However, a very straightforward yet resourceful tool has been overlooked in these efforts—the Lur’e system.
@@
Lur’e systems provide a flexible model for representing feedback systems with both system uncertainties and complicated nonlinearities like NNs, enabling the use of various stability analysis tools. This paper leverages the Lur’e framework to analyze the robustness of NN-controlled systems. We first explore the robustness of Lur’e systems under positivity constraints and structured uncertainties, then apply the framework to NN-controlled systems to enhance safety in applications.
@@
Lur’e systems have long served as a foundation for stability analysis of nonlinear systems, underpinning fundamental methods such as the Lur’e-Postnikov criterion, Aizerman and Kalman conjectures, the circle criterion, and their extensions. However, these classical approaches have limitations when dealing with system uncertainties, leaving room for further exploration. The existing literature on robustness analysis of Lur’e systems with structured uncertainties, which is the focus of our study, spans various approaches. For instance, studies such as [4, 5, 6] employ Popov criterion, $H_{\infty}$ control, and Lyapunov theory, respectively. Additionally, some works, like [7], explore more complex perturbation structures. There are recent works such as [8] that implicitly studied robust stability of Lur’e systems using Integral Quadratic Constraints (IQCs) and implemented it on an NN-controlled system. However, their proposed method suffers from high computational complexity when it comes to large systems. In contrast, this paper introduces a straightforward and scalable approach to robustness analysis of Lur’e systems. By utilizing positivity constraints on the system, we derive an explicit formula for the stability radius of Lur’e systems under a comprehensive class of structured uncertainties, improving upon prior works with similar uncertainty structure such as [9, 10] in both simplicity and computational efficiency.
@@
It is worth noting that positive-constrained systems are widely used in modeling numerous real-world systems, including engineering, biology, economics, and population dynamics. The positivity constraint offers a significant advantage in the stability and robustness analysis of linear systems [11, 12, 13]; however, its integration into nonlinear systems has not been extensively explored. In this paper, we exploit positivity constraints to develop efficient robustness analysis tools for nonlinear Lur’e systems under structured uncertainty. A key novelty of this paper is demonstrating how a linear system tool can be extended to analyze nonlinear robust stability.
@@
Lur’e systems have also been employed in the safety analysis of NNs. For instance, studies such as [14, 8, 15] implicitly utilize Lur’e systems alongside the S-procedure to define the forward reachable set of an NN and ensure the stability of NN-controlled systems. Our recent studies [16, 17] introduced a simple, scalable method for stability assurance in NN-controlled systems using the Lur’e framework. The studies [18, 19] implicitly utilize Lur’e system with reinforcement learning and iterative learning subject to time delays and nonlinear perturbations, respectively. Also, some studies [20] investigated the stability of feedback systems with ReLU NNs. However, none of them addressed a scalable and formally guaranteed robustness measure for NN feedback loops. To address this gap, our study proposes a straightforward and scalable method for calculating the stability radius of systems with NN feedback loops.
@@
In addition to robustness analysis, this study addresses sector bounds for NNs. Sector bounds are a classical tool in control theory, used to ensure stability in systems with nonlinearities by defining constraints within which the system’s nonlinearity must operate. NN bounds are also one of the leading topics in the machine learning community. There are extensive recent studies on NN bounding methods and refinement of the existing bounds. Examples are the evolution of bounding methods like CROWN [21], BnB [22], DeepPoly [23], Recurjac [24], and global sector bounds [16]. In this study, we build upon our results in Lur’e systems and propose a novel method for obtaining tight sector bounds for NNs.
@@ I-A Notation
The set of real numbers, $n$-dimensional real vectors, and $m\times n$-dimensional real matrices are denoted by $\mathbb{R}$, $\mathbb{R}^{n}$, and $\mathbb{R}^{m\times n}$, respectively. Additionally, $\mathbb{C}$ is set of complex numbers. The standard comparison operators <<, $\geq$, $\leq$, and >> are applied elementwise to vectors and matrices. $\mathbb{R}_{+}$ denotes the set of nonnegative real numbers. Moreover, $\mathbb{R}_{+}^{n}:=\{\nu\in\mathbb{R}^{n}:\nu\geq 0\}$ and $\mathbb{R}_{+}^{n\times m}:=\{A\in\mathbb{R}^{n\times m}:A\geq 0\}$ represents the set of vectors and matrices with real nonnegative elements. A Metzler matrix is characterized by having nonnegative off-diagonal entries. An M-matrix is defined as a matrix with non-positive off-diagonal elements and eigenvalues with non-negative real parts. A vector of zeros of size $n$ is denoted by $\mathbf{0}_{n}$. For a given matrix $A$, the order $|A|_{\text{abs}}$ means the elementwise absolute value of matrix $A$, and $\|A\|$ denotes a matrix norm of choice.
@@ II Proposed Theoretical Framework
In this section, we establish the theoretical frameworks that underpin our study, including the positive Lur’e system, positive Aizerman conjecture, and the stability radii of nonnegative and Metzler matrices.
@@ II-A Positive Lur’e Systems
Consider a linear time invariant (LTI) system of the form
$\dot{x}(t)=Ax(t)+Bu(t),\quad y(t)=Cx(t),$ (1)
@@
The system (1) is positive if and only if the matrix $A$ is a Metzler matrix, $B\in\mathbb{R}_{+}^{n\times m}$, and $C\in\mathbb{R}_{+}^{p\times n}$.
@@
A positive system of the form (1) is asymptotically stable if and only if there exists a positive vector $v$ such that $Av<0$.
Now, we officially define the positive Lur’e system. Consider a system formed by the interconnection of the linear system (1) and a static nonlinear feedback, represented as $u=\Phi(Cx,.)$. Fig. 1 depicts a schematic of this system. The overall dynamics of the system can be described by the following equation:
$\dot{x}=Ax(t)+B\Phi(Cx,t).$ (2)
@@
As part of the definition of Lur’e systems, it is necessary to establish sector bounds for the nonlinear component of the system. In the context of Multi Input Multi Output (MIMO) systems, the multivariable nature of $\Phi$ necessitates a generalization of the classical sector bound definition used in Single Input Single Output (SISO) Lur’e systems. For positive systems, the sector bound for the multivariable function $\Phi$ is best defined through componentwise inequalities. The function $\Phi$ is said to be sector bounded in $\Sigma_{1},\Sigma_{2}]$ if the following condition holds:
@@
where $\Sigma_{1},\Sigma_{2}\in\mathbb{R}^{m\times p}$, and $\Sigma_{1}\leq\Sigma_{2}$. The above formulation fully defines the MIMO positive Lur’e system under consideration. Based on this formulation, we proceed to introduce the positive Aizerman conjecture in the following subsection.
II-B Positive Aizerman Conjecture
The Aizerman conjecture is a classical method for absolute stability analysis of general Lur’e systems, originally formulated for SISO systems. While counterexamples exist in the general case, the conjecture holds under stricter conditions such as system positivity [26]. Furthermore, it can be extended to MIMO systems. By reducing nonlinear stability analysis to that of linear approximations, Aizerman conjecture offers a practical foundation for scalable and efficient stability analysis, particularly for systems with complex nonlinearities, such as NN-controlled systems.
The following theorem and remark confirm the validity of Aizerman conjecture for MIMO positive systems.
@@ II-B Positive Aizerman Conjecture
Consider the Lur’e system described by (2) where $B\geq 0$, $C\geq 0$, and the matrices $\Sigma_{1}$ and $\Sigma_{2}$ satisfy $\Sigma_{1}\leq\Sigma_{2}$ with appropriate dimensions. If the matrix $A+B\Sigma_{1}C$ is Metzler and $A+B\Sigma_{2}C$ is Hurwitz, then for any nonlinearity $\Phi$ within the sector $\Sigma_{1},\Sigma_{2}]$, as defined by (3), the Lur’e system (2) is globally exponentially stable.
@@
The Aizerman conjecture can be interpreted as a stability radius analysis of a linear system $\dot{x}=Ax$ subjected to an additive, structured perturbation given by
@@
The positive Aizerman conjecture asserts that if a system remains stable under all perturbations of a linear form $\{B\Sigma C:\quad\forall\Sigma\in[\Sigma_{1},\Sigma_{2}]\}$, then the system will also be stable for all corresponding nonlinear perturbations of the form $\{B\Phi(Cx):\forall\Phi\in[\Sigma_{1},\Sigma_{2}]\}$. In terms of stability radius, the positive Aizerman conjecture suggests that the ”real static nonlinear stability radius,” denoted $r_{\mathbb{R},\Phi}$, is equal to the ”real linear stability radius,” $r_{\mathbb{R}}$. This implies that the stability radius of a Lur’e system can be calculated by the stability radius of its linearizations.
@@ II-C Stability Radius of Positive Systems
where $A\in\mathbb{R}^{n\times n}$ is the system’s nominal matrix. $\Delta\in\mathbb{R}^{n\times n}$ represents a perturbation matrix and $\|\Delta\|$ is a norm representing the size of the perturbation. Suppose the following LTI system with structured perturbation:
@@
We focus on the case where $A$ is Metzler and stable. The stability assumption is included solely to guarantee the existence of the stability radius. Also, $D\geq 0$ and $E\geq 0$ are typical assumptions for systems with Metzler matrices and simplify the stability analysis.
The following theorem provides an explicit formula for the stability radius of Metzler matrices. For this class of matrices, it has been recognized that the real and complex stability radii are coinciding and given by a direct formula.
@@
Consider the linear system (5). Suppose that $A$ is Metzler and stable, $D\geq 0$ and $E\geq 0$. Let $\|\cdot\|$ be a chosen operator norm on $\mathbb{F}^{m\times p},\mathbb{F}=\mathbb{R},\mathbb{C}$. Then, the stability radii for system (5) under real perturbation $\Delta\in\mathbb{R}^{m\times p}$ $(r_{\mathbb{R}})$ or complex perturbations $\Delta\in\mathbb{C}^{m\times p}$ $(r_{\mathbb{C}})$ are given by
@@
Furthermore, if $\Delta$ is defined by the set $\Delta=\{S\odot\Delta:S_{ij}\geq 0\}$ with $\|\Delta\|=\max\{|\delta_{ij}|:\delta_{ij}\neq 0\}$ where $S\odot\Delta]_{ij}=S_{ij}\delta_{ij}$ represents the Schur product, then
@@ III Stability Radius of Positive Lur’e Systems
In this section, we formally define our problem and introduce our main result. We demonstrate and prove how we can utilize characteristics of Metzler matrices in linear systems for handling problems of nonlinear systems. Specifically, we study the stability radius of nonlinear Lur’e systems with structured perturbations under positivity constraints. Consider the positive Lur’e system described by (2), where the nonlinearity $\Phi(.)$ satisfies the sector condition (3) in $\Sigma_{1},\Sigma_{2}]$, with $\Sigma_{1},\Sigma_{2}\in\mathbb{R}^{m\times p}$ and $\Sigma_{1}\leq\Sigma_{2}$. We are interested in the robustness of this system to structured perturbations of the form:
@@
$\small r_{\mathbb{R}}(\Phi)$ $=\inf\{\|\Delta\|:\left(A+D\Delta E\right)x(t)$ (9)
$+B\Phi(Cx(t))\text{ is not globally asymptotically stable.}\}$
@@
Let $P,Q\in\mathbb{R}^{n\times n}$ be Metzler and Hurwitz stable matrices. Consider structured perturbations of the form:
@@
$\small r_{\mathbb{R}}(P)\;\leq\;r_{\mathbb{R}}(Q),$ (10)
where $r_{\mathbb{R}}(.)$ is calculated by (6) with $\|\cdot\|$ being a chosen operator norm that is monotone on the cone of nonnegative matrices (e.g., $\|\cdot\|_{1},\|\cdot\|_{\infty}$).
@@
Since $P$ is Metzler and Hurwitz, P-P has nonpositive off-diagonals and eigenvalues with positive real parts. This implies that P-P is an M-Matrix. The same applies to Q-Q.
Given $P\geq Q$, we have $P\leq-Q$. Therefore, Based on a classical property of nonsingular M-matrices
@@
Since we use a monotone operator norm $\|\cdot\|$, it follows that
@@
Taking reciprocals and removing the negative sign inside the norm yield the following inequality for the stability radii:
@@
We can also explain Lemma 3 in different words. It is well known that the interval positive systems $[Q,P]$ are Hurwitz stable if and only if $P$ is Hurwitz stable. This has been established in the theorem $4$ of reference [27] for positive discrete time systems, which is equally applicable for continuous-time positive systems, as it is shown in [28]. Now, without loss of generality, assume $D=E=I$. Since $P>Q$ and $P$ is asymptotically stable, it is closest to instability due to the continuity of stable eigenvalues[29, Lemma 3.5]. Thus, the distance of $P$ from the set of singular matrices is given by $\|P^{-1}\|^{-1}$, which is the smallest singular value of $P$ with respect to 22-norm. In other words, $r_{\mathbb{C}}=r_{\mathbb{R}}=\sigma_{\min}=\|P^{-1}\|^{-1},$ where $\sigma$ denotes the singular value of matrix $P$.
@@
Consider a Lur’e system described in (8). Given that $A+B\Sigma_{1}C$ is Metzler and $A+B\Sigma_{2}C$ is Hurwitz, the stability radius of positive Lur’e system $r_{\mathbb{R}}(\Phi)$ with respect to the structured perturbation $D\Delta E$ is given by
@@
Based on positive Aizerman conjecture and remark 1, the stability radius of the nonlinear Lur’e system can be measured by the stability radius of all the linear systems of the form $A+B\Sigma C,\Sigma\in[\Sigma_{1},\Sigma_{2}]$ where $\Phi$ is sector bounded in the same interval. In other words
$\small r_{\mathbb{R}}(\Phi)=\min_{\Sigma\in[\Sigma_{1},\Sigma_{2}]}r_{\mathbb{R}}(A+B\Sigma C).$ (12)
@@
$=r_{\mathbb{R}}(A+B\Sigma_{2}C).$ (13)
Now, (12) and (13) result in:
@@
Since the nonlinear function $\Phi$ lies in the sector $\Sigma_{1},\Sigma_{2}]$, and $A+B\Sigma_{2}C$ is Hurwitz, based on the implications of Remark 1, the same stability radius applies to the Lur’e system. ∎
It should be pointed out that the above theorem can also be stated in terms of other uncertainty structures, such as the one stated in Theorem 2 with the aid of equation (7).